OpenStudy (anonymous):

For f(x)= -sinx, x= pi/4 determine the equation for the tangent to the curve at the point with the given x-coordinate

OpenStudy (anonymous):

Substitute.\[f\left(\frac{\pi}{4}\right)=?\]

OpenStudy (anonymous):

the equation is y=-sin(x)

OpenStudy (anonymous):

? not quite sure what you mean by this question

OpenStudy (nottim):

maybe he was to look for the value of f(x) when x= pi/4?

OpenStudy (anonymous):

I need the derivative for f(x)= -sinx ... So that i can use y=mx+b to find the equation.

OpenStudy (anonymous):

What is the derivative of the sine function (something every calculus student must memorize!)?

OpenStudy (nottim):

isnt the derivative of -sin cosx?

OpenStudy (nottim):

oh shoot. i ruined that, srry yakey

OpenStudy (anonymous):

how'd you get cos(x) lol

OpenStudy (anonymous):

pull the negative out \[-\frac{d}{dx}[\sin(x)]\]

OpenStudy (nottim):

aww. wrogn again huh

OpenStudy (anonymous):

yakey you should remember all your derivatives. It will help for when you get to integration

OpenStudy (anonymous):

or anti derivatives.... basicaly this is how you learn